The coagulation of 100 mL of colloidal solution of gold is completely prevented by addition of 0.25 g of a substance ‘X’ to it, before adding 1 ml of 10% NaCl solution. The gold number of X is
    (A)    0.25
    (B)    25
    (C)    250
    (D)    2.5

Grade:upto college level

1 Answers

Radhika Batra
9 years ago

Amt. of protective colloid added in milligrams = 0.25 × 1000 = 250

    Hence By definition, the protective colloid required for 10 ml. =250/100×10=25